# Generative AI-enhanced Low-Altitude UAV-Mounted Stacked Intelligent Metasurfaces

### v1

@@ IV-C3 LBL-IPSO
Specially, for the $m$-th UAV-SIM, we optimize the phase shift $\bm{\Phi}^{l}_{m}$ of the $l$-th layer while keeping the phases of the other $(L-1)$ layers fixed. The channel gain $\mathbf{w}_{m}^{1\mathrm{H}}\mathbf{G}_{m}^{\mathrm{H}}\mathbf{h}_{m,k}$ can be expanded as follows:
@@
1 Let $\kappa=1$ and initialize $\bm{\Phi}_{m}^{1}$, $\bm{\Phi}_{m}^{2}$, …, $\bm{\Phi}_{m}^{L}$ randomly;
@@ IV-D Computational Complexity Analysis
Similarly, for ULOP, the total number of variables for $\mathrm{\mathbf{W}}$ and $\mathrm{\mathbf{A}}$ is $MK+2M$. The number of constraints in ULOP is $M(K-1)+M(M-1)/2$, and the number of iterations is $\sqrt{MK+2M}\log_{2}\big{(}{1}/{\varepsilon_{2}}\big{)}$, where $\varepsilon_{2}$ denotes the precision of CVX when solving ULOP. Thus, the computational complexity for ULOP is calculated as $\mathcal{O}[(MK)^{1.5}(MK+M^{2})\log_{2}{(1/\varepsilon_{2})}]$.
For USPSOP, when the HGPSO algorithm uses the LBL-IPSO strategy, calculating the $\mathbf{F}^{l\mathrm{H}}_{m}$ in line $5$ of Algorithm 2 requires $4(l-1)N^{2}$ floating-point computations, and calculating the $\mathbf{L}^{l\mathrm{H}}_{m}$ in line $6$ of Algorithm 2 requires $4(L-l)N^{2}$ floating-point computations. Therefore, the total computational complexity of LBL-IPSO is $\mathcal{O}[4N^{2}ML^{2}\kappa_{max}]$. Consequently, the total computational complexity for Algorithm 1 is $\mathcal{O}\{[(MK)^{2.5}\log_{2}(1/\varepsilon_{1})+(MK)^{1.5}(MK+M^{2})\log_{2}(1/\varepsilon_{2})+4N^{2}ML^{2}\kappa_{max}]\tau_{\max}\}$.
@@
It is evident that HGPSO using CVAE has lower computational complexity.
@@ V-A Simulation Setups
Regarding parameter configuration, we consider that $K=5$ ground users are distributed randomly in a square area with dimensions of $1000$ m ×$\times 1000$ m. Moreover, $3$ UAV-SIMs are flying in an aerial region at a fixed height $H=50$ m. The safety distance is established as $d_{\mathrm{min}}=100$ m, and the thickness of the SIM is established at $T_{\mathrm{SIM}}=5\lambda$ [6]. Therefore, for an $L$-layer SIM, the spacing between layers is $\delta=T_{\mathrm{SIM}}/L$. Additionally, we specify that the meta-atoms of the SIM are arranged in a square configuration, with each meta-atom having dimensions of $d_{x}d_{y}=(\lambda/2)^{2}$ [4].
[…]
@@ V-B Benchmarks
Figure 4: Network capacity $R$ versus SIM configuration. (a) Network capacity $R$ versus the number of metasurface layers $L$. (b) Network capacity $R$ versus the number of meta-atoms on each layer $N$.
@@ V-C Optimization Results
Fig. 3 illustrates the behavior of the loss function of the CVAE model over $300$ epochs with varying learning rates, where $L=4$ and $N=36$. As can be seen, the results demonstrate that all of the different learning rate configurations start around $23$. In particular, the learning rate of $41e-$ 4 performs best, with the lowest convergence value of around $11$, while the other four learning rates converge around $13$. Moreover, the $41e-$ 4 learning rate decreases rapidly in the first $50$ epochs and thereby converges at $100$ epochs. In contrast, the $21e-$ 2 learning rate shows a slower decline in loss values and requires approximately $150$ epochs to reach relative stability. In addition, learning rates of $31e-$ 3, $35e-$ 3 and $45e-$ 4 exhibit similar convergence patterns. However, the $45e-$ 4 learning rate shows a faster initial decline rate than the other two. Since the learning rate of $41e-$ 4 achieved the best results, we will use this learning rate for the CVAE model in subsequent simulations.
@@
Fig. 4 shows a comparison of network capacity performance across various algorithms at different SIM layer counts, where $N=36$. The CVAE and LBL-IPSO algorithms perform best across all layer counts. Between $1$ and $5$ layers, both methods exhibit a positive correlation with increasing SIM layer count, thereby reaching approximately $13$ bps at $L=5$. The performance declines slightly when the layer count becomes too large. Among these, CVAE performs slightly better than LBL-IPSO in layers $3$ to $5$. Although the RD algorithm performs well, it still lags behind CVAE and LBL-IPSO and exhibits significant variability, which is attributed to the random nature of RD. In the UD algorithm, performance increases from $3.8$ bps to $4.5$ bps as the number of layers increases, thus further indicating that increasing the number of SIM layers within a reasonable range can enhance the network capacity of the system. The swarm intelligence optimization algorithms PSO and DE perform relatively poorly among all comparison algorithms, which are significantly less effective than CVAE and LBL-IPSO. This phenomenon may be attributed to the large solution space, which consequently leads to suboptimal algorithm performance.
@@
Next, we explore the possible impact of different SIM configurations on network capacity from the perspective of the number of meta-atoms on each layer. Fig. 4 shows the performance of different algorithms in terms of network capacity as the number of meta-atoms varies, where $L=4$. The CVAE and LBL-IPSO algorithms demonstrate superior results in all configurations, with the network capacity of both algorithms being approximately $11$ bps when $N=16$. The network capacity increases as $N$ increases, thus reaching a maximum value of $13$ bps when $N=64$. This observation indicates that the multi-user interference cancellation capability of the SIM is positively correlated with the number of meta-atoms on each layer, which can be further verified through the trend of UD. The RD algorithm exhibits moderate performance with minor fluctuations, yet the same trend can still be observed. Furthermore, the network capacity under both PSO and DE can be steadily improved as the number of meta-atoms increases, thereby resulting in an enhancement of approximately $2$ bps for both algorithms. In comparison to Fig. 4, it is more advantageous to prioritize increasing the number of meta-atoms when modifying the configuration of the SIM.
@@
Fig. 5 illustrates the network capacity under different numbers of users within LAE networks, where $L=4$ and $N=36$. The network capacity under all the different algorithms decreases as the number of users increases, which is due to the constraint that a UAV-SIM can only serve one user, and a user can only receive one UAV-SIM service. Notably, CVAE and LBL-IPSO continue to be the algorithms with optimal performance, thereby yielding approximately $15$ bps with $3$ users, and the performance gradually decreases to approximately $9.4$ bps as the number of users increases to $8$. It should be noted that the network capacity decreases from $3.2$ bps initially to approximately $1.4$ bps in the absence of SIM service. When $K=3$, the network capacity with SIM is approximately $4.7$ times that without SIM, whereas it increases to approximately $5.4$ times when $K=6$ and to approximately $6.7$ times when $K=8$. This result thus demonstrates that, particularly within multi-user LAE deployment scenarios, the SIM can effectively mitigate the interference caused by multiple users.
@@
These results verify the superior performance of CVAE and LBL-IPSO. To further explore the time efficiency of different switching strategies in HGPSO, we compare and analyse their average execution time (AET). Under a unified CPU running environment with $N=36$, $30$ separate runs are performed for different SIM layers to calculate the average value and thus reduce the influence of outliers on the results. Fig. 6 shows the trend in AET for the CVAE model and the LBL-IPSO algorithm for different numbers of layers. As the complexity of the LAE scenario increases, it is clear that the two strategies show a significant difference in running efficiency. The AET for the LBL-IPSO algorithm starts at around $215$ seconds for a single layer SIM and increases with the number of layers, thereby reaching approximately $245$ seconds at $6$ layers, which is an increase of about $14\%$. In contrast, the CVAE model consistently maintains an AET between $222$ and $226$ seconds under the same conditions, which demonstrates greater stability with a smaller range of fluctuation. Compared with the LBL-IPSO algorithm, the AET is reduced by approximately $10\%$ at $6$ layers. This difference in trend suggests that the algorithmic efficiency of CVAE can effectively manage the computational burden of complex optimization problems in large-scale LAE network environments without time loss.

### v2

@@ IV-C3 LBL-IPSO
Specially, for the $m$-th UAV-SIM, we optimize the phase shift $\bm{\Phi}^{l}_{m}$ of the $l$-th layer while keeping the phases of the other (L1)(L-1) layers fixed. The channel gain $\mathbf{w}_{m}^{1\mathrm{H}}\mathbf{G}_{m}^{\mathrm{H}}\mathbf{h}_{m,k}$ can be expanded as follows:
@@
1Let $\kappa=1$ and initialize $\bm{\Phi}_{m}^{1}$, $\bm{\Phi}_{m}^{2}$, …, $\bm{\Phi}_{m}^{L}$ randomly;
@@ IV-D Computational Complexity Analysis
Similarly, for ULOP, the total number of variables for $\mathrm{\mathbf{W}}$ and $\mathrm{\mathbf{A}}$ is $MK+2M$. The number of constraints in ULOP is M(K1)+M(M1)/2M(K-1)+M(M-1)/2, and the number of iterations is $\sqrt{MK+2M}\log_{2}\big({1}/{\varepsilon_{2}}\big)$, where $\varepsilon_{2}$ denotes the precision of CVX when solving ULOP. Thus, the computational complexity for ULOP is calculated as $\mathcal{O}[(MK)^{1.5}(MK+M^{2})\log_{2}{(1/\varepsilon_{2})}]$.
For USPSOP, when the HGPSO algorithm uses the LBL-IPSO strategy, calculating the $\mathbf{F}^{l\mathrm{H}}_{m}$ in line $5$ of Algorithm 2 requires $4(l-1)N^{2}$ floating-point computations, and calculating the $\mathbf{L}^{l\mathrm{H}}_{m}$ in line $6$ of Algorithm 2 requires $4(L-l)N^{2}$ floating-point computations. Therefore, the total computational complexity of LBL-IPSO is $\mathcal{O}(4N^{2}ML^{2}\kappa_{max})$. Consequently, the total computational complexity for Algorithm 1 is $\mathcal{O}\{[(MK)^{2.5}\log_{2}(1/\varepsilon_{1})+(MK)^{1.5}(MK+M^{2})\log_{2}(1/\varepsilon_{2})+4N^{2}ML^{2}\kappa_{max}]\tau_{\max}\}$.
@@
Table II summarizes the computational complexity comparison of different approaches discussed in our work. It is evident that HGPSO using CVAE has lower computational complexity.
TABLE II: Computational complexity comparison of different algorithms.
Algorithm Time Complexity
Proposed HGPSO using LBL-IPSO $\mathcal{O}(4N^{2}ML^{2}\kappa_{\max})$
Proposed HGPSO using CVAE $\mathcal{O}(D_{n}z_{n}L_{n})$
@@ V-A Simulation Setups
Regarding parameter configuration, we consider that $K=5$ ground users are distributed randomly in a square area with dimensions of $1000$ m $\times 1000$ m. Moreover, $3$ UAV-SIMs are flying in an aerial region at a fixed height H=50H=50[49]. For the UAV energy consumption model parameters, we adopt representative values for a typical UAV platform as detailed in Table III [43]. The safety distance is established as $d_{\mathrm{min}}=100$ m, and the thickness of the SIM is established at $T_{\mathrm{SIM}}=5\lambda$ [6]. Therefore, for an $L$-layer SIM, the spacing between layers is $\delta=T_{\mathrm{SIM}}/L$. Additionally, we specify that the meta-atoms of the SIM are arranged in a square configuration, with each meta-atom having dimensions of $d_{x}d_{y}=(\lambda/2)^{2}$ [4].
We consider the radio frequency of $28$ GHz, and the corresponding wavelength is $\lambda=10.7$ mm [4]. Furthermore, the number of meta-atoms for each layer of the SIM is established at $N=36$. The noise power is assumed to be $\sigma^{2}_{m}=-110\text{ dBm},\forall m\in\mathcal{M}$, and the transmit power of each user is configured at $500\text{ mW}$. Moreover, the channel gain at a reference distance of 1 m is established at $\rho_{0}=(\lambda/4\pi)^{2}$ [6]. We establish the threshold $\epsilon$ for the AO strategy at $10^{-6}$, and the maximum number of iterations for AO is configured as $\tau_{\max}=50$. In addition, for Algorithm 2, we establish the number of iterations as $\kappa_{\max}=200$, thereby ensuring the quality of the solution. For the CVAE-M training, we utilize a dataset consisting of approximately 80,000 distinct entries collected from various simulation scenarios with comprehensive parameter variations including UAV-SIM positions, user positions, and channel environments to ensure dataset diversity and model generalization capability.
TABLE III: Notations and simulation values for UAV energy consumption model
Notation Physical meaning Simulation value
$W$ UAV-SIM weight (N) 120
$\rho$ Air density (kg/m3) 1.225
$R$ Rotor radius (m) 0.5
AA Rotor disc area, $A=\pi R^{2}$ (m2) 0.79
$\Omega$ Rotor angular velocity (rad/s) 400
$U_{tip}$ Tip speed of the rotor blade, $U_{tip}=\Omega R$ (m/s) 200
$s$ Rotor solidity 0.05
$\delta$ Profile drag coefficient 0.012
$d_{0}$ Equivalent flat plate area ratio 0.3
$P_{0}$ Profile power in hover (W) 580.7
$v_{0}$ Mean rotor induced velocity in hover (m/s) 7.87
$v_{m}$ UAV flying speed (m/s) 10.0
$P_{i}$ Induced power in hover (W) 944.9
@@ V-C Optimization Results
Figure 4: Network performance versus different loss functions.
Fig. 3 illustrates the behavior of the loss function of the CVAE model over $300$ epochs with varying learning rates, where $L=4$ and $N=36$. As can be seen, the results demonstrate that all of the different learning rate configurations start around $23$. In particular, the learning rate of 1e41e-4 performs best, with the lowest convergence value of around $11$, while the other four learning rates converge around $13$. Moreover, the 1e41e-4 learning rate decreases rapidly in the first $50$ epochs and thereby converges at $100$ epochs. In contrast, the 1e21e-2 learning rate shows a slower decline in loss values and requires approximately $150$ epochs to reach relative stability. In addition, learning rates of 1e31e-3, 5e35e-3 and 5e45e-4 exhibit similar convergence patterns. However, the 5e45e-4 learning rate shows a faster initial decline rate than the other two. Since the learning rate of 1e41e-4 achieved the best results, we will use this learning rate for the CVAE model in subsequent simulations.
@@
To further validate the effectiveness of our proposed approach, we conduct ablation studies demonstrating how different components of our loss function in Eq. (41) affect the overall performance. Specifically, we perform comparative simulations between the standard CVAE loss function and our improved formulation. The standard CVAE loss function focuses solely on reconstruction accuracy and latent space regularization, while our designed loss function incorporates an additional capacity-oriented term controlled by parameter $\beta_{3}$. The comparative results, as shown in Fig. 4, demonstrate significant performance advantages of our designed loss function, achieving approximately 400400% improvement in network capacity for single metasurface layer configurations compared to the standard approach. This performance gap persists across all configurations. These results validate that traditional CVAE require domain-specific modifications to excel in wireless optimization problems.
Figure 5: Network capacity $R$ versus SIM configuration. (a) Network capacity $R$ versus the number of metasurface layers $L$ with $K=5$ and $M=3$. (b) Network capacity $R$ versus the number of meta-atoms on each layer NN with $K=5$ and $M=3$. (c) Network capacity $R$ versus the number of metasurface layers $L$ with $K=8$ and $M=5$. (d) Network capacity $R$ versus the number of meta-atoms on each layer NN with $K=8$ and $M=5$.
@@
Fig. 5 shows a comparison of network capacity performance across various algorithms at different SIM layer counts, where $N=36$. The CVAE and LBL-IPSO algorithms perform best across all layer counts. Between $1$ and $5$ layers, both methods exhibit a positive correlation with increasing SIM layer count, thereby reaching approximately $13$ bps at $L=5$. The performance declines slightly when the layer count becomes too large. Among these, CVAE performs slightly better than LBL-IPSO in layers $3$ to $5$. Although the RD algorithm performs well, it still lags behind CVAE and LBL-IPSO and exhibits significant variability, which is attributed to the random nature of RD. In the UD algorithm, performance increases from $3.8$ bps to $4.5$ bps as the number of layers increases, thus further indicating that increasing the number of SIM layers within a reasonable range can enhance the network capacity of the system. The swarm intelligence optimization algorithms PSO and DE perform relatively poorly among all comparison algorithms, which are significantly less effective than CVAE and LBL-IPSO. This phenomenon may be attributed to the large solution space, which consequently leads to suboptimal algorithm performance. To evaluate scalability, Fig. 5 shows similar trends in an expanded scenario with $K=8$ users and $M=5$ UAV-SIMs, where our CVAE model continues to outperform all benchmark methods and reaches nearly $12$ bps with $6$ layers, representing a $14.3\%$ improvement over LBL-IPSO. This demonstrates the effective scalability of our proposed approach across varying network complexities.
@@
Next, we explore the possible impact of different SIM configurations on network capacity from the perspective of the number of meta-atoms on each layer. Fig. 5 shows the performance of different algorithms in terms of network capacity as the number of meta-atoms varies, where $L=4$. The CVAE and LBL-IPSO algorithms demonstrate superior results in all configurations, with the network capacity of both algorithms being approximately $11$ bps when $N=16$. The network capacity increases as $N$ increases, thus reaching a maximum value of $13$ bps when $N=64$. This observation indicates that the multi-user interference cancellation capability of the SIM is positively correlated with the number of meta-atoms on each layer, which can be further verified through the trend of UD. The RD algorithm exhibits moderate performance with minor fluctuations, yet the same trend can still be observed. Furthermore, the network capacity under both PSO and DE can be steadily improved as the number of meta-atoms increases, thereby resulting in an enhancement of approximately $2$ bps for both algorithms. In comparison to Fig. 5, it is more advantageous to prioritize increasing the number of meta-atoms when modifying the configuration of the SIM. Similarly, Fig. 5 demonstrates the scalability of our approach in the larger network scenario, where the CVAE model achieves approximately $11$ bps with $64$ meta-atoms, which confirms the robust performance across both network scaling and metasurface dimensionality.
@@
Fig. 6 illustrates the network capacity under different numbers of users within LAE networks, where $L=4$ and $N=36$. The network capacity under all the different algorithms decreases as the number of users increases, which is due to the constraint that a UAV-SIM can only serve one user, and a user can only receive one UAV-SIM service. Notably, CVAE and LBL-IPSO continue to be the algorithms with optimal performance, thereby yielding approximately $15$ bps with $3$ users, and the performance gradually decreases to approximately $9.4$ bps as the number of users increases to $8$. It should be noted that the network capacity decreases from $3.2$ bps initially to approximately $1.4$ bps in the absence of SIM service. When $K=3$, the network capacity with SIM is approximately $4.7$ times that without SIM, whereas it increases to approximately $5.4$ times when $K=6$ and to approximately $6.7$ times when $K=8$. This result thus demonstrates that, particularly within multi-user LAE deployment scenarios, the SIM can effectively mitigate the interference caused by multiple users. These substantial performance gains stem from the multi-layer SIM architecture that provides significantly more degrees of freedom for wavefront control and enables superior interference mitigation compared to conventional single-layer designs.
@@
These results verify the superior performance of CVAE and LBL-IPSO. To further explore the time efficiency of different switching strategies in HGPSO, we compare and analyse their average execution time (AET). Under a unified CPU running environment with $N=36$, $30$ separate runs are performed for different SIM layers to calculate the average value and thus reduce the influence of outliers on the results. Fig. 7 shows the trend in AET for the CVAE model and the LBL-IPSO algorithm for different numbers of layers. As the complexity of the LAE scenario increases, it is clear that the two strategies show a significant difference in running efficiency. The AET for the LBL-IPSO algorithm starts at around $215$ seconds for a single layer SIM and increases with the number of layers, thereby reaching approximately $245$ seconds at $6$ layers, which is an increase of about $14\%$. In contrast, the CVAE model consistently maintains an AET between $222$ and $226$ seconds under the same conditions, which demonstrates greater stability with a smaller range of fluctuation. Compared with the LBL-IPSO algorithm, the AET is reduced by approximately $10\%$ at $6$ layers. This difference in trend suggests that the algorithmic efficiency of CVAE can effectively manage the computational burden of complex optimization problems in large-scale LAE network environments without time loss. This is because traditional optimization methods face exponentially growing computational burdens with more SIM layers and meta-atoms, while our CVAE approach shifts the computational burden to an offline training phase and enables rapid online inference. Consequently, this makes CVAE the preferred choice for real-time UAV-SIM deployments in dynamic environments where rapid adaptation is crucial.